OpenStudy (anonymous):

3x+2y=-9 -10x+5y=-5 solve using procces of elimation

OpenStudy (cwrw238):

start by multiplying equation 1 by 5 and equation 2 by 5 then subtract to eliminate y

OpenStudy (anonymous):

so in the first equation do i multiply by 5?

OpenStudy (cwrw238):

yes - dont forget multiply every term by 5

OpenStudy (anonymous):

Ok thank you =D

OpenStudy (cwrw238):

give it a go

OpenStudy (asnaseer):

you need to multiply equation 1 by some number (call it A) and equation 2 by some other number (call it B) such that either the coefficients of x or y become the same.

OpenStudy (asnaseer):

so for x (for example) we need 3A = 10B or for y, we need 2A = 5B

OpenStudy (asnaseer):

can you think of what A and B should B to make either of these two equal?

OpenStudy (anonymous):

No

OpenStudy (asnaseer):

ok, lets say we want to eliminate y, then we need to find two numbers A and B such that: 2A = 5B we could pick A=5 and B=2 to get: 2*5 = 5*2

OpenStudy (asnaseer):

so now if we multiply the first equation by 5 and the second one by 2 we get: 5(3x + 2y) = 5(-9) ==> 15x + 10y = -45 2(-10x + 5y) = 2(-5) ==> -20x + 10y = -10

OpenStudy (asnaseer):

now we have two equations where the coefficients of y are the same, so we can subtract one from the other to eliminate y. do you know how to do that?

OpenStudy (asnaseer):

15x + 10y = -45 - equation 3 -20x + 10y = -10 - equation 4 equation 3 minus equation 4 will give us: (15 - (-20))x = -45 - (-10) (15 + 20)x = -45 + 10 35x = -35 x = -1 you can now substitute this back into either your first or second equation to find y.